OpenStudy (anonymous):

I'll medal and fan! Which ratio is equivalent to 4 : 28? A. 1 : 8 B. 1 : 7 C. 8 : 52 D. 3 : 14

OpenStudy (anonymous):

4 : 28 = 4 / 28 = (4 ÷ 4) / (28 ÷ 4) = 1 / 7 = 1 : 7 (A)

OpenStudy (anonymous):

thanks